OpenStudy (anonymous):

how would you graph |x| + |y| = 1? And why is it not a function?

OpenStudy (anonymous):

do you know what makes a graph a function?

OpenStudy (tkhunny):

It's not a function because it isn't. Vertical line test? For x>0 and y > 0 (1st Quadrant), this is x + y = 1 For x<0 and y > 0 (2md Quadrant), this is -x + y = 1 For x<0 and y < 0 (3rd Quadrant), this is -x - y = 1 You do the last one - Quadrant 4. Graph the four lines and see what gives!

OpenStudy (anonymous):

what he said is exactly what I was going to say haha

OpenStudy (anonymous):

and don't forget about your absolute value signs

OpenStudy (anonymous):

you got this or no? it is not nearly as hard as it looks

OpenStudy (anonymous):

Quadrant 4: x>0 y<0 x - y =1 I got it! thanks everyone:)